# SECap: Speech Emotion Captioning with Large Language Model

### 16 Dec 2023 v1

@@ Abstract
Speech emotions are crucial in human communication and are extensively used in fields like speech synthesis and natural language understanding. Most prior studies, such as speech emotion recognition, have categorized speech emotions into a fixed set of classes. Yet, emotions expressed in human speech are often complex, and categorizing them into predefined groups can be insufficient to adequately represent speech emotions. On the contrary, describing speech emotions directly by means of natural language may be a more effective approach. Regrettably, there are not many studies available that have focused on this direction. Therefore, this paper proposes a speech emotion captioning framework named SECap, aimed at effectively describing speech emotions using natural language. Owing to the impressive capabilities of large language models in language comprehension and text generation, SECap employs LLAMA as the text decoder to allow the production of coherent speech emotion captions. In addition, SECap leverages HuBERT as the audio encoder to extract general speech features and Q-Former as the Bridge-Net to provide LLAMA with emotion-related speech features. To accomplish this, Q-Former utilizes mutual information learning to disentangle emotion-related speech features and speech contents, while implementing contrastive learning to extract more emotion-related speech features. The results of objective and subjective evaluations demonstrate that: 1) the SECap framework outperforms the HTSAT-BART baseline in all objective evaluations; 2) SECap can generate high-quality speech emotion captions that attain performance on par with human annotators in subjective mean opinion score tests.
@@ Introduction
However, traditional SER exhibits limitations, as single-word labels often lack nuance, failing to convey detailed emotional information like intensity and fluctuations. Speech emotions are typically multifaceted, encompassing diverse affective states (e.g., simultaneous happiness and nervousness) within one utterance. Classifying speech into a single emotion category may inadequately capture authentic emotion. Additionally, the inherently subjective perception of emotions leads to potential variability in emotion classification among individuals interpreting complicated speech. Considering the limitations of speech emotion classification, employing natural language sentences rather than labels could be a promising strategy to describe speech emotions more precisely. Motivated by the recent progress of the automated audio captioning (AAC) task (Han et al. 2021; Chen et al. 2020; Ye et al. 2021) which employs natural language to describe acoustic events in audio, we present the Speech Emotion Captioning (SEC) task and propose an innovative SECap framework, comprising an audio encoder, a Bridge-Net, and a text decoder, to characterize human speech emotions using natural language. To our knowledge, this is among the pioneering works in this direction.
@@
Figure 1: Comparison of Speech Emotion Recognition (SER) model and the proposed SECap, the SER model generates emotion labels, while SECap generates natural language emotion descriptions derived from the same speech.
@@
Large Language Models (LLMs) (Singhal et al. 2023; Kojima et al. 2022; Black et al. 2022) have revolutionized natural language understanding, impacting areas like machine translation. By analyzing vast textual data, LLMs learn linguistic patterns and generate natural prose. Open-source models like BLOOM (Scao et al. 2022) and ChatGLM (Du et al. 2022) have fostered growth in the LLM community. Also, researchers explore LLMs’ performance in multimodal interactions, aspiring for models capable of managing audio, vision, and text modalities, reflecting human daily interactions. The first approach uses LLMs as task orchestrators, connecting downstream models like AudioGPT (Huang et al. 2023) and HuggingGPT (Shen et al. 2023) for specialized tasks. The second approach positions LLMs as multitask processors, mapping modal tasks to a unified space. For example, BLIP-2 maps images to text space using Q-Former, while Video-LLAMA (Zhang, Li, and Bing 2023) maps audio and vision modalities via Q-Former.
@@
Inspired by the AAC task, SECap employs an encoder-decoder architecture, as illustrated in Figure 1. The audio encoder extracts speech features, and Bridge-Net extracts emotion-related speech features and transforms them into the text decoder’s feature space. The text decoder then generates speech emotion captions based on these features.
@@
The HuBERT is to derive speech embedding for its powerful speech feature extraction capability. However, frame-level HuBERT features can lead to heavy computation cost. We employ Q-Former-based Bridge-Net to compress features. Meanwhile, acoustic information is more directly associated with speech emotion, while content information is obtainable from transcriptions. Thus, the Bridge-Net is used to extract emotion-related acoustic information and eliminate content information. We employ LLAMA as the text decoder for generating speech emotion captions, leveraging its exceptional language comprehension capabilities. Aligning with LLAMA’s input format, we position L-Embedding between the ·“BOS” and a prompt. This method constrains LLAMA’s output space via the prompt, yielding more accurate speech emotion captions.
@@
Owing to the redundancy of HuBERT speech features, the Q-Former is designed to compress and extract emotion-related speech features and consists of self-attention, cross-attention, and linear layers. Q-queries are learnable parameters for extracting speech embedding.
@@
where $W_{q_{\text{self}}}\in\mathbb{R}^{d_{q}\times d_{k}}$, $W_{k_{\text{self}}}\in\mathbb{R}^{d_{q}\times d_{k}}$, and $W_{v_{\text{self}}}\in\mathbb{R}^{d_{q}\times d_{v}}$ are the learnable weight matrices for queries, keys, and values in the self-attention mechanism, and $d_{k}$ and $d_{v}$ are the dimensions of keys and values. The output of the self-attention mechanism $A_{\text{self}}$ is then used as queries for the cross-attention mechanism, while the speech embedding $S\in\mathbb{R}^{n_{s}\times T_{s}\times d_{s}}$ serve as keys and values:
@@
where $A_{\text{cross}}\in\mathbb{R}^{n_{s}\times n_{q}\times d_{q}}$, $W_{q}\in\mathbb{R}^{d_{q}\times d_{k}}$, $W_{k}\in\mathbb{R}^{d_{s}\times d_{k}}$, and $W_{v}\in\mathbb{R}^{d_{s}\times d_{v}}$ are the learnable weight matrices for queries, keys, and values in the cross-attention mechanism.
This approach enables the attention mechanism to retrieve features related to Q-queries within the speech embedding. The Q-Embedding $Q_{e}\in\mathbb{R}^{n_{s}\times d_{q}}$, which is the output of the Q-Former, exhibits a fixed length that remains unchanged regardless of the input speech length, resulting in enhanced generalization performance for different speech lengths.
@@
To provide LLAMA with more content-unrelated and emotion-related speech features, we simultaneously incorporate both the human-labeled speech emotion captions and the transcriptions. As depicted in Figure 3, these are passed through a Q-Former that is largely consistent with the original, except for the absence of the cross-attention module. This process yields T-Embedding $Q_{t}\in\mathbb{R}^{n_{s}\times T_{t}\times d_{q}}$ and C-Embedding $Q_{c}\in\mathbb{R}^{n_{s}\times T_{c}\times d_{q}}$, where $T_{c}$ and $T_{t}$ denote the length of the caption and the transcription, respectively. we employ Speech-Transcription Mutual Information Learning to disentangle speech features from speech content. Additionally, Speech-Caption Contrastive Learning is utilized to extract more emotion-related speech features.
@@
Considering potential speech content factors that could interfere with emotion assessment, such as expressing joyous statements calmly, this content may affect emotion assessment. To reduce the correlation between speech features and content, mitigating the speech content’s impact on LLAMA’s speech emotion caption generation, we propose Speech-Transcription Mutual Information Learning. As illustrated in Figure 3, we introduce both Speech Embedding and Trans Embedding into the Q-Former simultaneously, yielding $Q_{e}$ and T-Embedding $Q_{t}$. This method enables the comparison of speech and its content within a unified representation space. In order to evaluate the correlation between $Q_{e}$ and $Q_{t}$, we adopt mutual information as the metric. The mutual information $I(Q_{t};Q_{e})$ is defined as:
@@
However, due to the undetermined specific distribution and high-dimensional nature of $Q_{e}$ and $Q_{t}$, we cannot directly calculate their mutual information. While prior methods such as MINE (Belghazi et al. 2018) and infoNCE (Van Den Oord, Vinyals et al. 2017) can estimate the lower bound of mutual information, our primary objective is to reduce the mutual information between $Q_{e}$ and $Q_{t}$. Consequently, we utilize the vCLUB (Cheng et al. 2020) method to approximate the upper bound of the mutual information. Specifically, we denote the mutual information obtained by the vCLUB method as $\mathcal{U}(Q_{t};Q_{e})$, which is defined as:
@@
where $n$ is the number of samples, and $x_{i}$ and $y_{i}$ represent the $i$-th samples of $Q_{t}$ and $Q_{e}$, respectively.
Our primary objective is to reduce the mutual information between $Q_{e}$ and $Q_{t}$. Consequently, we use the mutual information upper bound $\mathcal{U}(Q_{t};Q_{e})$ as the loss function.
@@
Due to the high dimensionality and redundancy in speech representations, speech features contain abundant information such as content and background noise, with only a fraction being emotion-related. To alleviate the complexity of processing speech features by LLAMA, we aim for Q-Former to extract features highly correlated with speech emotion caption features, consequently bridging the gap between speech features and text modality. As illustrated in Figure 3, our objective is to minimize the distance between $Q_{e}$ and $Q_{c}$, prompting Q-Former to extract more emotion-related features and progressively approach the text modality. Drawing inspiration from CLAP (Wu et al. 2023), we employ a contrastive learning approach to accurately represent distances between $Q_{e}$ of distinct speech samples, ensuring that speech with similar emotions yield closer $Q_{e}$ distances, while those with dissimilar emotions result in farther $Q_{e}$ distances.
To mitigate the influence of similar emotions in negative samples during contrastive learning, we partition EMOSpeech into N distinct categories based on human-labeled speech emotion labels. This approach guarantees substantial differences in speech emotion captions across categories, enhancing the model’s discriminative capacity throughout the learning process. During each training step, we select K speech-caption pairs from each $N$ set, ensuring that for each $Q_{e}$, there is 1 corresponding $Q_{c}$ (denoted as $d_{i}$),$(K-1)$ $Q_{c}$ with similar emotions (denoted as $p_{i}$), and $(NK-K)$ $Q_{c}$ with dissimilar emotions (denoted as $u_{i}$).
We opt to use cosine similarity $\mathcal{S}$ to measure the distance between $Q_{e}$ and $Q_{t}$. For enhanced comparative learning, we design the training method as follows:
@@
To enhance speech emotion caption generation with LLAMA, we devise a two-stage training process. The first stage compresses HuBERT-extracted speech features to obtain emotion-relevant attributes, while the subsequent stage aligns these features with LLAMA’s representation space. Inspired by Blip-2, we initialize the Q-Former using pre-trained parameters from $\text{BERT}_{\text{base}}$ (Devlin et al. 2019).
In the first training stage, we combine STMIL and SCCL for collaborative training. Specifically, our training loss for the first phase is as follows:
@@
where the weighting coefficients $w_{T1}$ and $w_{T2}$ control the contribution of STMIL and SCCL in the first training stage.
To optimally integrate Q-Former-extracted speech features into LLAMA, we fine-tune the Q-Former and the projection layer during the second training stage, maintaining LLAMA’s frozen parameters. We also insert a “BOS” token before the L-Embedding to align with the inference format. To improve SECap’s generalization ability, we devise 30 semantically akin sentences, each instructing to “portray the speaker’s emotion in a single Chinese sentence.” During training, we randomly choose a sentence to concatenate after L-Embedding. Subsequently, We append the human-labeled speech caption $C_{1}$ after the prompt and employ the teacher-forcing approach to enable LLAMA to generate the caption $C_{2}$. Cross-entropy loss (CELoss) guides Q-Former and the projection layer learning, as illustrated below:
$\mathcal{L}_{T2}=\text{CELoss}(C_{1},C_{2})$ (7)
@@
Due to the lack of publicly available SEC datasets, we utilized an internal dataset called EMOSpeech. EMOSpeech dataset consists of 5 female and 2 male speakers, totaling 41.6 hours of speech covering 30526 sentences, sampled at a rate of 24kHz. Each speech in EMOSpeech has three to five human-labeled speech emotion captions and human-labeled speech emotion labels provided by different annotators, along with its corresponding transcription.
Owing to the subjectivity of speech emotion captions, we developed a three-step process to minimize variations among annotators [^2] EMOSpeech is currently a Mandarin Speech Emotion Caption dataset since most of the annotators are native Chinese speakers. and enhance caption quality. First, annotators identify the overall emotion in a single word, serving as the speech emotion label in EMOSpeech, without considering speech content. Next, they indicate the emotion’s specific intensity (e.g., “strongly” or “slightly”). Finally, based on prior steps, annotators provide a comprehensive natural language caption for the speech emotion.
Upon constructing the EMOSpeech dataset, we randomly select 600 sentences for testing [^3] The test set is publicly available at project’s GitHub repository., 600 sentences for validation, and the remaining 29,326 sentences for training.
@@
Since there was no existing method to evaluate speech emotion captions, we devised both objective and subjective evaluation methods based on the nature of the SEC task.
@@
Our experiments are conducted exclusively on the EMOSpeech dataset. We choose the HuBERT-large model, pre-trained on the 10k-hour WenetSpeech (Zhang et al. 2022) L subset, as the audio encoder. Due to the original LLAMA’s limited proficiency in understanding Chinese, we choose an enhanced version of LLAMA (Cui, Yang, and Yao 2023) finetuned with Chinese datasets as the text decoder.
@@
Table 1 illustrates the number of trainable and total parameters during the two stages of the training process.
Table 2: Objective experiment results on the performance analysis of SECap. Specifically, “Raw Trans” denotes the raw transcription of the speech.“Q-Emb” denotes the Q-Embedding.“T-Emb” denotes the T-Embedding. “—” refers to not included. Higher scores mean better performance.
@@
This experiment aims to demonstrate the effectiveness of SECap. Given speech content’s impact on emotion, we incorporate transcriptions as additional input and design multiple comparison groups for thorough analysis. Specifically, when incorporating transcriptions, we either use the raw transcriptions processed through the LLAMA tokenizer or apply the T-Embedding processed through a projection layer as an extra input, which is then concatenated between BOS Emb and L-Embedding. Subsequently, we append the Prompt Emb and feed it into LLAMA. In some of the comparison groups, speech features are not used, the former processed transcriptions are directly concatenated between BOS Emb and Prompt Emb. To thoroughly assess the experiment, we employ a blend of objective and subjective evaluation metrics, as outlined in the Evaluation Metric chapter.
@@
Table 2 demonstrates that most metrics show consistent trends across diverse experiments; thus, our analysis will chiefly concentrate on insights derived from the two Chinese sentence similarity models.
@@
In the subjective experiment, we randomly select a test set of 50 sentences. We apply all the experimental methods listed in Table 2 to generate corresponding speech emotion captions. To enable a more comprehensive comparison, we also include human-labeled speech emotion labels (Human Label) and emotion categories (SER Model Label) identified by a competitive pre-trained Chinese SER model from HuggingFace [^4] https://huggingface.co/xmj2002/hubert-base-ch-speech-emotion-recognition. We request evaluators to score the nine types of texts according to the evaluation details provided by the supervisor in the Evaluation Metrics. In total, 15 evaluators participate in the subjective tests, with the results presented in Figure 4.
Figure 4: Subjective experiment results on the performance analysis of SECap.
@@
Table 3: Experimental results on the ablation study of different model components
@@
This experiment seeks to investigate the impact of various training methods for the Q-Former on generating speech emotion captions. While maintaining the audio encoder as HuBERT and the text decoder as LLAMA, we conducted a series of comparisons, including whether to use STMIL or SCCL in the first training stage, and whether to freeze the Q-Former in the second training stage.
Table 4: Experimental results on the effect of different methods of training the Q-Former

### 19 Dec 2023 v2

@@ Abstract
Speech emotions are crucial in human communication and are extensively used in fields like speech synthesis and natural language understanding. Most prior studies, such as speech emotion recognition, have categorized speech emotions into a fixed set of classes. Yet, emotions expressed in human speech are often complex, and categorizing them into predefined groups can be insufficient to adequately represent speech emotions. On the contrary, describing speech emotions directly by means of natural language may be a more effective approach. Regrettably, there are not many studies available that have focused on this direction. Therefore, this paper proposes a speech emotion captioning framework named SECap, aiming at effectively describing speech emotions using natural language. Owing to the impressive capabilities of large language models in language comprehension and text generation, SECap employs LLaMA as the text decoder to allow the production of coherent speech emotion captions. In addition, SECap leverages HuBERT as the audio encoder to extract general speech features and Q-Former as the Bridge-Net to provide LLaMA with emotion-related speech features. To accomplish this, Q-Former utilizes mutual information learning to disentangle emotion-related speech features and speech contents, while implementing contrastive learning to extract more emotion-related speech features. The results of objective and subjective evaluations demonstrate that: 1) the SECap framework outperforms the HTSAT-BART baseline in all objective evaluations; 2) SECap can generate high-quality speech emotion captions that attain performance on par with human annotators in subjective mean opinion score tests.
@@ Introduction
However, traditional SER exhibits limitations, because single-word labels often lack nuances, failing to convey detailed emotional information like intensity and fluctuations. Speech emotions are typically multifaceted, encompassing diverse affective states (e.g., simultaneous happiness and nervousness) within one utterance. Classifying speech into a single emotion category may inadequately capture authentic emotion. Additionally, the inherently subjective perception of emotions leads to potential variability in emotion classification among individuals interpreting complicated speech. Considering the limitations of speech emotion classification, employing natural language sentences rather than labels could be a promising strategy to describe speech emotions more precisely. Motivated by the recent progress of the Automated Audio Captioning (AAC) task (Han et al. 2021; Chen et al. 2020; Ye et al. 2021) which employs natural language to describe acoustic events in audio, we present the Speech Emotion Captioning (SEC) task and propose an innovative SECap framework, comprising an audio encoder, a Bridge-Net, and a text decoder, to characterize human speech emotions using natural language. To our knowledge, this is among the pioneering works in this direction.
@@
Figure 1: Comparison of Speech Emotion Recognition (SER) model and the proposed SECap. The SER model generates emotion labels, while SECap generates natural language emotion descriptions derived from the speech.
@@
Large Language Models (LLMs) (Singhal et al. 2023; Kojima et al. 2022; Black et al. 2022) have revolutionized natural language understanding. By analyzing vast textual data, LLMs learn linguistic patterns and generate natural prose. Open-source models like BLOOM (Scao et al. 2022) and ChatGLM (Du et al. 2022) have fostered growth in the LLM community. Also, researchers explore LLMs’ performance in multimodal interactions, aspiring for models capable of managing audio, vision, and text modalities, reflecting human daily interactions. The first approach uses LLMs as task orchestrators, connecting downstream models like AudioGPT (Huang et al. 2023) and HuggingGPT (Shen et al. 2023) for specialized tasks. The second approach positions LLMs as multitask processors, mapping modal tasks to a unified space. For example, BLIP-2 maps images to text space using Q-Former, while Video-LLaMA (Zhang, Li, and Bing 2023) maps audio and vision modalities via Q-Former.
@@
Inspired by the AAC task, SECap employs encoder-decoder architecture, as illustrated in Figure 1. The audio encoder extracts speech features, and Bridge-Net extracts emotion-related speech features and transforms them into the text decoder’s feature space. The text decoder then generates speech emotion captions based on these features.
@@
The HuBERT is to derive speech embedding for its powerful speech feature extraction capability. However, frame-level HuBERT features can lead to heavy computation cost. We employ Q-Former-based Bridge-Net to compress features. Meanwhile, acoustic information is more directly associated with speech emotion, while content information is obtainable from transcriptions. Thus, the Bridge-Net is used to extract emotion-related acoustic information and eliminate content information. We employ LLaMA as the text decoder for generating speech emotion captions, leveraging its exceptional language comprehension capabilities. Aligning with LLaMA’s input format, we position L-Embedding between the “BOS” and a prompt. This method constrains LLaMA’s output space via the prompt, yielding more accurate speech emotion captions.
@@
Owing to the redundancy of HuBERT speech features, Q-Former is designed and adopted to compress and extract emotion-related speech features which consists of self-attention, cross-attention, and linear layers. Q-queries are learnable parameters for extracting speech embedding.
@@
where $W_{q_{\text{self}}}\in\mathbb{R}^{d_{q}\times d_{k}}$, $W_{k_{\text{self}}}\in\mathbb{R}^{d_{s}\times d_{k}}$, and $W_{v_{\text{self}}}\in\mathbb{R}^{d_{q}\times d_{v}}$ are the learnable weight matrices for queries, keys, and values in the self-attention mechanism, and $d_{k}$ and $d_{v}$ are the dimensions of keys and values. The output of the self-attention mechanism $A_{\text{self}}\in\mathbb{R}^{n_{q}\times d_{v}}$ are then used as the queries for the cross-attention mechanism, while the speech embedding $S\in\mathbb{R}^{n_{s}\times T_{s}\times d_{s}}$ serve as keys and values:
@@
where $A_{\text{cross}}\in\mathbb{R}^{n_{s}\times n_{q}\times d_{v}}$ represents the cross-attention output, while $W_{q}\in\mathbb{R}^{d_{v}\times d_{k}}$, $W_{k}\in\mathbb{R}^{d_{s}\times d_{k}}$, and $W_{v}\in\mathbb{R}^{d_{s}\times d_{v}}$ are the learnable weight matrices for queries, keys, and values in the cross-attention mechanism.
This approach enables the attention mechanism to retrieve features related to Q-queries within the speech embedding. Specifically, the output of the Q-Former, denoted as the Q-Embedding $Q_{e}\in\mathbb{R}^{n_{s}\times d_{q}}$, maintains a fixed length that is independent of the length of the input speech. This fixed-length representation leads to improved generalization performance across speech inputs of varying lengths.
@@
To provide LLaMA with more content-unrelated and emotion-related speech features, we simultaneously incorporate both the human-labeled speech emotion captions and the transcriptions. As depicted in Figure 3, these are passed through a Q-Former that is largely consistent with the original, except for the absence of the cross-attention module. This process yields C-Embedding $Q_{c}\in\mathbb{R}^{n_{s}\times T_{c}\times d_{q}}$ and T-Embedding $Q_{t}\in\mathbb{R}^{n_{s}\times T_{t}\times d_{q}}$, where $T_{c}$ and $T_{t}$ denote the length of the caption and the transcription, respectively. We employ Speech-Transcription Mutual Information Learning to disentangle speech features from speech content. Additionally, Speech-Caption Contrastive Learning is utilized to extract more emotion-related speech features.
@@
The speech content can potentially impact emotion assessment, for example, expressing joyous statements calmly. To minimize the correlation between speech features and content, thereby mitigating the speech content’s impact on LLaMA’s speech emotion caption generation, we propose Speech-Transcription Mutual Information Learning. As illustrated in Figure 3, we introduce both Speech Embedding and Trans Embedding into the Q-Former simultaneously, yielding Q-Embedding $Q_{e}$ and T-Embedding $Q_{t}$. This enables the comparison of speech and its content within a unified representation space. To evaluate the correlation between $Q_{e}$ and $Q_{t}$, we adopt mutual information $I(Q_{t};Q_{e})$ as the metric:
@@
However, direct computation of mutual information between $Q_{e}$ and $Q_{t}$ is infeasible due to their unknown, high-dimensional nature. While prior methods such as MINE (Belghazi et al. 2018) and infoNCE (Van Den Oord, Vinyals et al. 2017) can estimate the lower bound of mutual information, they are not suitable for controlling the minimization process. Following vCLUB (Cheng et al. 2020), we use Equation (4) to estimate the upper bound of mutual information and employ it as a loss function to reduce the correlation between speech features and content.
@@
The equation includes conditional probabilities $q(y_{i}|x_{i})$ and $q(y_{j}|x_{i})$, representing the probabilities of the $i$-th and $j$-th $Q_{e}$ samples given the $i$-th $Q_{t}$ sample. The logarithm captures the dissimilarity between $Q_{e}$ conditioned on $Q_{t}$, and summing over all pairwise combinations provides the upper bound mutual information measure between $Q_{e}$ and $Q_{t}$.
@@
Due to the high dimensionality and redundancy in speech representations, speech features contain abundant information such as content and background noise, with only a fraction being emotion-related. To alleviate the complexity of processing speech features by LLaMA, we aim for Q-Former to extract features highly correlated with speech emotion caption, consequently bridging the gap between speech features and text modality. As illustrated in Figure 3, our objective is to minimize the distance between $Q_{e}$ and C-Embedding $Q_{c}$, prompting Q-Former to extract more emotion-related features and progressively approach the text modality. Drawing inspiration from CLAP (Wu et al. 2023), we employ a contrastive learning approach to accurately represent distances between $Q_{e}$ of distinct speech samples, ensuring that speech with similar emotions yield closer $Q_{e}$ distances, while those with dissimilar emotions result in farther $Q_{e}$ distances.
To mitigate the influence of similar emotions in negative samples during contrastive learning, we partition the dataset into $N$ distinct categories based on human-labeled speech emotion labels. This guarantees substantial differences in speech emotion captions across categories, thereby enhancing the model’s discriminative capacity throughout the learning process. During each training step, we select $K$ speech-caption pairs from each of the $N$ sets, ensuring that for each $Q_{e}$ (referred to as $e_{i}$), there is 1 corresponding $Q_{c}$ (referred to as $d_{i}$),$(K-1)$ $Q_{c}$ with similar emotions (referred to as $p_{i}$), and $(NK-K)$ $Q_{c}$ with dissimilar emotions (referred to as $u_{i}$).
We opt to use cosine similarity $\mathcal{S}$ to measure the distance between $Q_{e}$ and $Q_{c}$. For enhanced contrastive learning, we design the training method as follows:
@@
To enhance speech emotion caption generation with LLaMA, we devise a two-stage training process. The first stage compresses HuBERT-extracted speech features to obtain emotion-relevant attributes, while the subsequent stage aligns these features with LLaMA’s representation space.
In the first training stage, we combine STMIL and SCCL for collaborative training as in Figure 3, while keeping the HuBERT model frozen. Inspired by BLIP-2, we initialize the Q-Former using pre-trained parameters from $\text{BERT}_{\text{base}}$ (Devlin et al. 2019). Specifically, the training loss is:
@@
where the weighting coefficients $w_{T1}$ and $w_{T2}$ control the contribution of STMIL and SCCL.
In the second training stage, we fine-tune the Q-Former and the projection layer to effectively integrate Q-Former-extracted speech features into LLaMA. Meanwhile, the parameters of LLaMA and HuBERT remain frozen. We insert a “BOS” token before the L-Embedding to align with the inference format. To improve SECap’s generalization ability, we devise 30 semantically akin sentences, each instructing to “portray the speaker’s emotion in a single Chinese sentence.” During training, we randomly choose a sentence to concatenate after the L-Embedding. Subsequently, we append the human-labeled speech caption $C$ after the prompt and employ the teacher-forcing approach to enable LLaMA generate caption $\hat{C}$. Cross-entropy loss (CELoss) is then adopted as the training objective:
$\mathcal{L}_{T2}=\text{CELoss}(C,\hat{C})$ (7)
@@
Due to the lack of publicly available SEC datasets, we utilize an internal dataset called EMOSpeech. EMOSpeech dataset consists of 5 female and 2 male speakers, totaling 41.6 hours of speech covering 30526 sentences, sampled at a rate of 24kHz. Each speech in EMOSpeech has three to five human-labeled speech emotion captions and human-labeled speech emotion labels provided by different annotators, along with its corresponding transcription.
As for labeling, we begin with 50 sample audio clips for independent annotator labeling and hold a discussion session for annotators to review annotations and establish standardized rules based on collective input. The annotation process has three levels: identifying overall emotion using a single word, describing emotion intensity, and providing a comprehensive sentence considering emotion, volume, and speech rate. With these guidelines, annotators consistently label the dataset. To ensure annotation quality, we conduct consistency checks by randomly selecting 5 out of every 100 clips for review by other annotators, upholding high standards throughout the dataset construction.
Upon constructing the EMOSpeech dataset, we randomly select 600 sentences for testing, 600 sentences for validation, and the remaining 29,326 sentences for training [^2] Please refer to project’s GitHub repository for detailed dataset construction process, where the test set is also publicly available..
@@
Since there is no existing method to evaluate speech emotion captions, we devise both objective and subjective evaluation methods based on the nature of the SEC task.
@@
Our experiments are conducted exclusively on the EMOSpeech dataset. We choose the HuBERT-large model, pre-trained on the 10k-hour WenetSpeech (Zhang et al. 2022) L subset, as the audio encoder. Due to the original LLaMA’s limited proficiency in understanding Chinese, we choose an enhanced version of LLaMA (Cui, Yang, and Yao 2023) finetuned with Chinese datasets as the text decoder [^3] Specific experimental details are given at GitHub repository..
@@
Table 2: Objective experiment results on performance analysis. “Raw Trans” denotes raw transcription of speech.“Q-Emb” denotes Q-Embedding.“T-Emb” denotes T-Embedding. “—” refers to not included. Higher scores mean better performance.
@@
This experiment aims to demonstrate the effectiveness of SECap. Given speech content’s impact on emotion, we incorporate transcriptions as additional input and design multiple comparison groups for thorough analysis. Specifically, when incorporating transcriptions, we either use the raw transcriptions processed through the LLaMA tokenizer or apply the T-Embedding processed through a projection layer as an extra input, which is then concatenated between BOS Emb and L-Embedding. Subsequently, we append the Prompt Emb and feed it into LLaMA. In some of the comparison groups, speech features are not used, the former processed transcriptions are directly concatenated between BOS Emb and Prompt Emb.
@@
Table 2 illustrates consistent trends in most metrics across various experiments. Therefore, our analysis will primarily focus on the insights obtained from the two Chinese sentence similarity models.
@@
In the subjective experiment, we randomly select a test set of 50 sentences. We apply all the methods listed in Table 2 to generate corresponding speech emotion captions. To enable a more comprehensive comparison, we also include human-labeled speech emotion labels (Human Label) and emotion categories (SER Model Label) identified by a competitive pre-trained Chinese SER model from HuggingFace [^4] https://huggingface.co/xmj2002/hubert-base-ch-speech-emotion-recognition. We request evaluators to score the nine types of texts according to the Subjective Evaluation details provided in Evaluation Metric section. 15 evaluators participate in the tests, with the results presented in Figure 4.
Figure 4: Subjective experiment results on performance analysis
@@
Table 3: Experiment results on the ablation study of different model components
@@
This experiment seeks to investigate the impact of various training methods for the Q-Former on generating speech emotion captions. While maintaining the audio encoder as HuBERT and the text decoder as LLaMA, we conduct a series of comparisons, including whether to use STMIL or SCCL in the first training stage, and whether to freeze the Q-Former in the second training stage.
Table 4: Experiment results on the effect of different methods of training the Q-Former